Find the cube root of each of the following numbers by prime factorisation method.

 

(ii) 512

 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(ii) 512

Solution:

512 = 2×2×2×2×2×2×2×2×2

By grouping the factors in triplets of equal factors, 512 = (2×2×2)×(2×2×2)×(2×2×2)

Here, 512 can be grouped into triplets of equal factors,

∴ 512 = 2×2×2 = 8

Hence, 8 is cube root of 512.




Thanks